In a class of 45 students, 27 are democrats, 8 are business majors and 6 of the business majors are democrats. If one student is randomly selected from the class,find the probability of selecting a democrat or a business major.
Answer by fractalier(6550)

P(demo or bus) = P(demo) + P(bus) - P(demo and bus)
P(demo or bus) = 27/45 + 8/45 - 6/45 = 29/45